# Differentiable Discrete Elastic Rods for Real-Time Modeling of Deformable Linear Objects

### v1

@@ Abstract
This paper addresses the task of modeling Deformable Linear Objects (DLOs), such as ropes and cables, during dynamic motion over long time horizons. This task presents significant challenges due to the complex dynamics of DLOs. To address these challenges, this paper proposes differentiable Discrete Elastic Rods For deformable linear Objects with Real-time Modeling (DEFORM), a novel framework that combines a differentiable physics-based model with a learning framework to model DLOs accurately and in real-time. The performance of DEFORM is evaluated in an experimental setup involving two industrial robots and a variety of sensors. A comprehensive series of experiments demonstrate the efficacy of DEFORM in terms of accuracy, computational speed, and generalizability when compared to state-of-the-art alternatives. To further demonstrate the utility of DEFORM, this paper integrates it into a perception pipeline and illustrates its superior performance when compared to the state-of-the-art methods while tracking a DLO even in the presence of occlusions. Finally, this paper illustrates the superior performance of DEFORM when compared to state-of-the-art methods when it is applied to perform autonomous planning and control of DLOs.
@@ 1 Introduction
Robotic manipulation tasks such as surgical suturing or vehicle wire harness assembly [1, 2, 3] are challenging problems because they require accurately and dynamically manipulating objects such as ropes or cables over long time horizons ($>1s$[4]. These objects are often categorized as Deformable Linear Objects (DLOs) because they have complex dynamics due to bending and twisting deformations. These nonlinear dynamics make manipulating DLOs challenging because computationally expensive models are required to accurately predict their behavior. Further, practical applications, such as vehicle wire harness assembly, often result in DLO occlusions, meaning perception systems cannot robustly estimate the entire DLO configuration. For robotic manipulators to successfully manipulate DLOs in practical applications, a novel modeling method capable of quick and accurate predictions dynamic DLO behavior, even in the presence of occlusions, is needed.
@@ 2 Related Work
Physics-Based Modeling. Traditionally, DLOs have been modeled using physics-based approaches. Examples include mass-spring system models [16, 17, 18], Position-Based Dynamics (PBD) models [19, 20, 21], Discrete Elastic Rod (DER) models [13, 14, 22, 23], or Finite Element Method (FEM) models [24, 25, 26]. Although easy to implement, the mass-spring system imposes unnecessary stiffness to enforce inextensibility, resulting in instability during simulation. FEM models accurately modeling DLO behavior, but are impractical for real-time prediction due to their prohibitive computational costs. PBD and DER methods have been applied in planning and control frameworks for manipulation tasks due to their computational efficiency [20, 21, 27, 28, 29]. However, PBD models are sensitive to parameter selection and assume quasi-static behavior. As illustrated in this paper, this results in low accuracy during dynamic manipulation. Though DER models have been qualitatively validated to describe the 3D behavior of DLOs, their quantitative performance while describing dynamic motion is lacking as we illustrate in this paper.
Differentiable Modeling Framework. Differentiable models are computational frameworks where every operation is differentiable with respect to the input parameters. Importantly, a differentiable model enables integration into a learning framework to improve learning and can enable efficient parameter auto-tuning via gradient-based optimization during training. There are differentiable models for deformable objects, such as DiffCloth [16, 17] and XPBD [21], but these models tend to suffer from numerical instability.This makes it challenging to apply these modeling frameworks while trying to perform parameter identification for real-world DLO manipulation. To address these limitations, this paper proposes a differentiable DER model, called DDER, which is incorporated into a learning framework that achieves numerical stability when predicting dynamic DLO behavior. This results in better predictive performance when compared to state-of-the-art methods and enables auto-tuning of DLO parameters using limited real-world data.
@@ 3 Preliminaries
Deformable Linear Object (DLO) Model. To model a DLO, we separate the DLO into an indexed set of $n$ vertices at each time instance. The ground truth locations of vertex $i$ at time $t$ is denoted by $\textbf{x}^{\text{i}}_{\text{t}}\in\mathbb{R}^{3}$ and the set of all $n$ vertices is denoted by $\textbf{X}_{\text{t}}$. Let $\textit{{M}}^{\text{i}}\in\mathbb{R}^{3\times 3}$ denote the mass matrix of vertex $i$. Our objective is to model the dynamic behavior of a DLO that is being controlled by a robot with following assumption: the first and last edge of the DLO are each held by a distinct end effector. Note this assumption is commonly made during DLO manipulation [30, 20, 19, 7, 31]. Using this assumption, let the orientation and position of the two end effectors at time $t$ be denoted by $\textbf{u}_{t}\in\mathbb{R}^{12}$, which we refer to as the input. The velocity of the vertices of the DLO can be denoted as $\textbf{V}_{\text{t}}=(\textbf{X}_{\text{t}}-\textbf{X}_{\text{t-1}})/\Delta\text{t}$. Finally, note that we distinguish between ground truth elements and predicted elements by using the circumflex symbol (e.g., $\textbf{X}_{t}$ is the ground truth set of vertices at time $t$ and $\hat{\textbf{X}}_{\text{t}}$ is the predicted set of vertices at time $t$).
Modeling DLO with DER for Dual Manipulation. DLOs primarily exhibit three distinct modes: bending, twisting, and stretching [13]. This paper focuses on DLOs that are inextensible (e.g., cables and wires), and do not experience stretching. DER models introduce a single scalar $\theta^{\text{i}}_{t}\in\mathbb{R}$ to represent bending and twisting in $\mathbb{R}^{3}$. This parameter describes the potential energy $P(\textbf{X}_{t},\bm{\theta}_{t}(\textbf{X}_{t}),\bm{\alpha})$ of the DLO that arises due to bending or twisting, where $\bm{\alpha}$ denotes the vector of the material properties of each vertex of the DLO (i.e., their mass, bending modulus, and twisting modulus) [13, Section 4.2]. Notably, DER models assume that the DLO reaches an equilibrium state between each simulation time step to obtain $\bm{\theta}^{*}_{t}$:
@@ 4.2 Integration Method with Residual Learning
where DNN denotes a deep neural network. We utilize a structure similar to residual learning [36], with one major difference: we define a shortcut connection that is equal to the integration method that is derived from the DDER physics-based modeling. Unlike pure learning methods, the short cut connection grounds the DNN in physical laws while also leveraging data-driven insights. As shown in the ablation study results, this approach significantly improves modeling performance. Implementation details of the DNN structure can be found in Appendix A.1. Note that this contribution can also be applied to standard DER models, but does not improve performance as much as it does when combined with our DDER modeling approach as we illustrate in this paper.
@@ 4.3 Improved Inextensiblity Enforcement with PBD
where $\textbf{r}^{i}$ is the vector between $\hat{\textbf{x}}^{i}_{t+1}$ and a common rotation center. Note in particular that if the mass matrix associated with each vertex is not identical, then the correction proposed in (11) leads to a violation of (12) . To address this issue, we modify the correction proposed in (11) by multiplying it by a $\beta\in\mathbb{R}$ that is inversely correlated with the weight ratio between successive vertices (i.e., let $\beta^{i}=\frac{\textit{{M}}^{i+1}}{(||\textit{{M}}^{i}||_{2}+||\textit{{M}}^{i+1}||_{2})}$ and $\beta^{i+1}=\frac{\textit{{M}}^{i}}{(||\textit{{M}}^{i}||_{2}+||\textit{{M}}^{i+1}||_{2})}$). Then, if we let the correction between successive steps be equal to $\beta^{i}\Delta\hat{\textbf{x}}^{i}_{t+1}$, one can ensure that (12) is satisfied. Pseudo code to enforce inextensibility while preserving momentum can be found in Appendix A.2.
@@ 5 Experiments and Results
This section covers the real-world evaluation of DEFORM, including: 1) a quantitative comparison of modeling accuracy with state-of-the-art methods, 2) computational speed assessment, 3) an ablation study on parameter auto-tuning, residual learning, and multi-step training, 4) using DEFORM to track DLOs with an RGB-D camera, and 5) a shape matching manipulation demonstration. Code and data will be provided upon final paper acceptance. Additional information is in Appendix B.
@@ 5.2 Modeling Results
Ablation Study. Results are summarized in Table 2 and show the effect of each DEFORM contribution. Note that a standard DER method is included as a baseline and does not include any of the DEFORM contributions. The results of the study show that residual learning has the biggest improvement on the baseline performance, but the auto-tuning of material properties has a non-trivial affect as well. Further, the improved inextensibility enforcement makes a large improvement for more compliant DLOs, such as DLO1. This is because the more compliant DLOs swing more during dynamic behavior, so ensuring momentum is conserved greatly improves prediction accuracy. The results for DLO4 and DLO5, as well as an additional ablation study, are in Appendix B.3 and B.4, respectively. Note that the stiffnesses of the tested DLOs can be found in Appendix Table 4.
@@
State Estimation. We next evaluate the performance of DEFORM when it is used in concert with sensor observations for state estimation. These experiments included occlusions of the DLO that arose from the manipulators or humans moving in front of the sensor. In this experiment, we compare how well the state estimation framework described in Appendix C does when using the DEFORM model versus the Bi-LSTM model. Bi-LSTM was chosen because it was the second best model in terms of model accuracy. We also compare how well each of these methods do with varying frequencies of sensor measurement updates. As in the earlier experiment, we evaluate the L1 norm of the prediction over a 5s long prediction horizon. Experimental results are shown in Figure 4. Unsurprisingly, each method improves when given access to more frequent sensor updates. This illustrates that the state estimation algorithm described in Appendix C.2 behaves appropriately. Notably, the DEFORM model with no state estimation does better than the Bi-LSTM model even when the Bi-LSTM model is given access to 1 fps updates. The method using the DEFORM model with access to 30 fps sensor measurements is the most effective.
@@
This experiment uses ARMOUR, a receding-horizon trajectory planner and tracking controller framework, to manipulate DLOs. ARMOUR uses either DEFORM or Bi-LSTM to predict the manipulated DLO state and tries to match the final DLO configuration with a desired configuration. Details of combining ARMOUR with the DLO simulators are in Appendix B.5. We compare the performance of each method on 100 random cases in a PyBullet simulation and 20 random cases in the real world. If the Euclidean distance between each individual pair of ground truth and desired DLO vertices is less than 0.05 meters, then the trial is a success. ARMOUR is given 30 seconds to complete the task, otherwise the trial is marked as a failure. The results are summarized in Table 3 and Figure 5 shows one of the real-world results. A video of these experiments can be found in the supplementary material.
@@ 6 Conclusion and Limitations
This paper presents DEFORM, a method that embeds residual learning in a novel differentiable Deformable Linear Object (DLO) simulator. This enables accurate modeling and real-time prediction of dynamic DLO behavior over long time horizons during dynamic manipulation. To demonstrate DEFORM’s efficacy, this paper performs a comprehensive set of experiments that evaluate accuracy, computational speed, and its usefulness for manipulation and perception tasks. When compared to the state-of-the-art, DEFORM is more accurate while being sample efficient without sacrificing computational speed. This paper also illustrates the practicality of applying DEFORM for state estimation to track a DLO during dynamic manipulation in environments with occlusions. Furthermore, DEFORM is combined with a manipulation planning and control framework for 3D shape matching of a DLO in both simulations and real-world scenarios, achieving a higher task success rate compared to the closest-performing baseline.
@@
Appendix A Algorithmic Details
A.1 DNN Structure
@@
A.2 Summary and Discussion of Improved Inextensiblity Enforcement with PBD

### v3

@@ Abstract
This paper addresses the task of modeling Deformable Linear Objects (DLOs), such as ropes and cables, during dynamic motion over long time horizons. This task presents significant challenges due to the complex dynamics of DLOs. To address these challenges, this paper proposes differentiable Discrete Elastic Rods For deformable linear Objects with Real-time Modeling (DEFORM), a novel framework that combines a differentiable physics-based model with a learning framework to model DLOs accurately in real-time. The performance of DEFORM is evaluated in an experimental setup involving two industrial robots and a variety of sensors. A comprehensive series of experiments demonstrate the efficacy of DEFORM in terms of accuracy, computational speed, and generalizability when compared to state-of-the-art alternatives. To further demonstrate the utility of DEFORM, this paper integrates it into a perception pipeline and illustrates its superior performance when compared to the state-of-the-art methods while tracking a DLO even in the presence of occlusions. Finally, this paper illustrates the superior performance of DEFORM when compared to state-of-the-art methods when it is applied to perform autonomous planning and control of DLOs.
@@ 1 Introduction
Robotic manipulation tasks such as surgical suturing or vehicle wire harness assembly [1, 2, 3] are challenging problems because they require accurately and dynamically manipulating DLOs over long time horizons ($>1s$[4]. Nonlinear dynamics make manipulating DLOs challenging because computationally expensive models are required for accurate prediction. Further, practical applications such as vehicle wire harness assembly often result in DLO occlusions, meaning perception systems cannot robustly estimate the entire DLO configuration. For robotic manipulators to successfully manipulate DLOs, a novel modeling method capable of quick and accurate predictions of dynamic DLO behavior, even in the presence of occlusions, is needed.
@@ 2 Related Work
Physics-Based Modeling. Traditionally, DLOs have been modeled using physics-based approaches. Examples include mass-spring system models [16, 17, 18], Position-Based Dynamics (PBD) models [19, 20, 21], Discrete Elastic Rod (DER) models [13, 14, 22, 23], or Finite Element Method (FEM) models [24, 25, 26]. Although easy to implement, the mass-spring system imposes unnecessary stiffness to enforce inextensibility, resulting in instability during simulation. FEM models accurately modeling DLO behavior, but are impractical for real-time prediction due to their prohibitive computational costs. PBD and DER methods have been applied in planning and control frameworks for manipulation tasks due to their computational efficiency [20, 21, 27, 28, 29]. However, PBD models are sensitive to parameter selection and assume quasi-static behavior. As illustrated in this paper, this results in low accuracy during dynamic manipulation. Though DER models have been qualitatively validated to describe the 3D behavior of DLOs, we illustrate that their quantitative performance while describing dynamic motion is lacking.
Differentiable Modeling Framework. Differentiable models are computational frameworks where every operation is differentiable with respect to the input parameters. Importantly, a differentiable model enables integration into a learning framework to improve learning and can enable efficient parameter auto-tuning via gradient-based optimization during training. There are differentiable models for deformable objects, such as DiffCloth [16, 17] and XPBD [21], but these models tend to suffer from numerical instability. This makes it challenging to apply these modeling frameworks while trying to perform parameter identification for real-world DLO manipulation. To address these limitations, this paper proposes a differentiable DER model, called DDER, which is incorporated into a learning framework to achieve numerical stability. This results in better predictive performance for dynamic DLO behavior when compared to state-of-the-art methods and enables auto-tuning of DLO parameters using limited real-world data.
@@ 3 Preliminaries
Deformable Linear Object (DLO) Model. To model a DLO, we separate the DLO into an indexed set of $n$ vertices at each time instance. The ground truth locations of vertex $i$ at time $t$ is denoted by $\textbf{x}^{\text{i}}_{\text{t}}\in\mathbb{R}^{3}$ and the set of all $n$ vertices is denoted by $\textbf{X}_{\text{t}}$. Let $\textit{{M}}^{\text{i}}\in\mathbb{R}^{3\times 3}$ denote the mass matrix of vertex $i$. Our objective is to model the dynamic behavior of a DLO that is being controlled by a robot with following assumption: the first and last edge of the DLO are each held by a distinct end effector. Note this assumption is commonly made during DLO manipulation [30, 20, 19, 7, 31]. Using this assumption, let the orientation and position of the two end effectors at time $t$ be denoted by $\textbf{u}_{t}\in\mathbb{R}^{12}$, which we refer to as the input. The velocity of the vertices of the DLO can be denoted as $\textbf{V}_{\text{t}}=(\textbf{X}_{\text{t}}-\textbf{X}_{\text{t-1}})/\Delta\text{t}$. Note that $\textbf{V}_{\text{t}}$ is an approximation of the actual velocity. Finally, note that we distinguish between ground truth elements and predicted elements by using the circumflex symbol (e.g., $\textbf{X}_{t}$ is the ground truth set of vertices at time $t$ and $\hat{\textbf{X}}_{\text{t}}$ is the predicted set of vertices at time $t$).
Modeling DLOs with DER for Dual Manipulation. DLOs primarily exhibit three distinct deformation modes: bending, twisting, and stretching [13]. This paper focuses on DLOs that are inextensible (e.g., cables and wires), and therefore do not experience stretching. DER models introduce a single scalar $\theta^{\text{i}}_{t}\in\mathbb{R}$ to represent bending and twisting in $\mathbb{R}^{3}$. This parameter describes the potential energy $P(\textbf{X}_{t},\textbf{u}_{t},\bm{\theta}_{t},\bm{\alpha})$ of the DLO that arises due to bending or twisting, where $\bm{\alpha}$ denotes the vector of the material properties of each vertex of the DLO (i.e., their mass, bending modulus, and twisting modulus) [13, Section 4.2]. Notably, DER models assume that the DLO reaches an equilibrium state between each simulation time step to obtain $\bm{\theta}^{*}_{t}$:
@@ 4.2 Integration Method with Residual Learning
where DNN denotes a deep neural network. We utilize a structure similar to residual learning [36], with one major difference: we define a shortcut connection that is equal to the integration method that is derived from the DDER physics-based modeling. Unlike pure learning methods, the short cut connection grounds the DNN in physical laws while also leveraging data-driven insights. As shown in the ablation study results, this approach significantly improves modeling performance. Additioanl discussion of section can be found in Appendix B.5. Implementation details of the DNN structure can be found in Appendix B.1. Note that this contribution can also be applied to standard DER models, but does not improve performance as much as it does when combined with our DDER modeling approach as we illustrate in this paper.
@@ 4.3 Improved Inextensiblity Enforcement with PBD
where $\textbf{r}^{i}$ is the vector between $\hat{\textbf{x}}^{i}_{t+1}$ and a common rotation center. Note in particular that if the mass matrix associated with each vertex is not identical, then the correction proposed in (11) leads to a violation of (12) . To address this issue, we modify the correction proposed in (11) by multiplying it by a $\beta\in\mathbb{R}$ that is inversely correlated with the weight ratio between successive vertices (i.e., let $\beta^{i}=\frac{\textit{{M}}^{i+1}}{(||\textit{{M}}^{i}||_{2}+||\textit{{M}}^{i+1}||_{2})}$ and $\beta^{i+1}=\frac{\textit{{M}}^{i}}{(||\textit{{M}}^{i}||_{2}+||\textit{{M}}^{i+1}||_{2})}$). Then, if we let the correction between successive steps be equal to $\beta^{i}\Delta\hat{\textbf{x}}^{i}_{t+1}$, one can ensure that (12) is satisfied. Verification proof can be found in Appendix B.2. Pseudo code to enforce inextensibility while preserving momentum can be found in Appendix B.3.
@@ 5 Experiments and Results
This section covers the real-world evaluation of DEFORM, including: 1) a quantitative comparison of modeling accuracy with state-of-the-art methods, 2) computational speed assessment, 3) an ablation study on parameter auto-tuning, residual learning, and multi-step training, 4) using DEFORM to track DLOs with an RGB-D camera, and 5) a shape matching manipulation demonstration. Code and data will be provided upon final paper acceptance. Additional information is in Appendix C.
@@ 5.2 Modeling Results
Ablation Study. Results are summarized in Table 2 and show the effect of each DEFORM contribution. Note that a standard DER method is included as a baseline and does not include any of the DEFORM contributions. The results of the study show that residual learning has the biggest improvement on the baseline performance, but the auto-tuning of material properties has a non-trivial affect as well. Further, the improved inextensibility enforcement makes a large improvement for more compliant DLOs, such as DLO1. This is because the more compliant DLOs swing more during dynamic behavior, so ensuring momentum is conserved greatly improves prediction accuracy. The results for DLO4 and DLO5, as well as an additional ablation study, are in Appendix C.3 and C.4, respectively. Note that the stiffnesses of the tested DLOs can be found in Appendix Table 4.
State Estimation. We next evaluate the performance of DEFORM when it is used in concert with
@@
sensor observations for state estimation. These experiments included occlusions of the DLO that arose from the manipulators or humans moving in front of the sensor. In this experiment, we compare how well the state estimation framework described in Appendix D does when using the DEFORM model versus the Bi-LSTM model. Bi-LSTM was chosen because it was the second best model in terms of model accuracy. We also compare how well each of these methods do with varying frequencies of sensor measurement updates. As in the earlier experiment, we evaluate the L1 norm of the prediction over a 5s long prediction horizon. Experimental results are shown in Figure 4. Unsurprisingly, each method improves when given access to more frequent sensor updates. This illustrates that the state estimation algorithm described in Appendix D.2 behaves appropriately. Notably, the DEFORM model with no state estimation does better than the Bi-LSTM model even when the Bi-LSTM model is given access to 1 fps updates. The method using the DEFORM model with access to 30 fps sensor measurements is the most effective.
@@
This experiment uses ARMOUR [37], a receding-horizon trajectory planner and tracking controller framework, to manipulate DLOs. ARMOUR uses either DEFORM or Bi-LSTM to predict the manipulated DLO state and tries to match the final DLO configuration with a desired configuration. Details of combining ARMOUR with the DLO simulators are in Appendix C.6. We compare the performance of each method on 100 random cases in a PyBullet simulation and 20 random cases in the real world. If the Euclidean distance between each individual pair of ground truth and desired DLO vertices is less than 0.05 meters, then the trial is a success. ARMOUR is given 30 seconds to complete the task, otherwise the trial is marked as a failure. The results are summarized in Table 3 and Figure 5 shows one of the real-world results. A video of these experiments can be found in the supplementary material.
@@ 6 Conclusion and Limitations
This paper presents DEFORM, a method that embeds residual learning in a novel differentiable Deformable Linear Object (DLO) simulator. This enables accurate modeling and real-time prediction of dynamic DLO behavior over long time horizons during dynamic manipulation. To demonstrate DEFORM’s efficacy, this paper performs a comprehensive set of experiments that evaluate accuracy, computational speed, and its usefulness for manipulation and perception tasks. When compared to the state-of-the-art, DEFORM is more accurate while being sample efficient without sacrificing computational speed. This paper also illustrates the practicality of applying DEFORM for state estimation to track an occluded DLO during dynamic manipulation. Furthermore, DEFORM is combined with a manipulation planning and control framework for 3D shape matching of a DLO in both simulations and real-world scenarios, achieving a higher task success rate compared to the closest-performing baseline.
@@
Acknowledgments
The authors would like to gratefully thank reviewers for giving useful comments. This work is supported by the Ford Motor Company and the Air Force Office of Scientific Research under the Award No: MURI FA9550-23-1-0400.
Appendix A Additional DER Background
Figure 6: An illustration of DER coordinate frames.
Bishop Frame. To accurately approximate the potential energy of the DLO that arises due to bending or twisting, DER first assigns Bishop frames $\{\mathbf{t}^{i},\mathbf{u}^{i},\mathbf{v}^{i}\}$ as reference frame onto edge $i$. $\mathbf{t}^{i}\in\mathbb{R}^{3}$ represents the unit tangent vector along edge $i$. The vector $\mathbf{u}^{i}\in\mathbb{R}^{3}$ is defined iteratively through a rotation operation $\mathbf{u}^{i}=\mathbf{R}^{i}\cdot\mathbf{u}^{i-1}$. This rotation is governed by a rotation matrix $\mathbf{R}^{i}\in\mathbb{R}^{3\times 3}$, which satisfies following conditions: $\mathbf{t}^{i}=\mathbf{R}^{i}\cdot\mathbf{t}^{i-1}$ and $\mathbf{t}^{i-1}\times\mathbf{t}^{i}=\mathbf{R}^{i}\cdot(\mathbf{t}^{i-1}\times\mathbf{t}^{i})$. $\mathbf{v}^{i}$, orthogonal to $\mathbf{t}^{i}$ and $\mathbf{u}^{i}$, is defined by $\mathbf{v}^{i}=\mathbf{t}^{i}\times\mathbf{u}^{i}$. This transportation results in the least amount of twist, making it ideal for adapting the Bishop frame as the most natural reference frame. Further details of the Bishop frame can be found in [40].
Material Frame. DER further introduces a scalar $\theta^{i}$ to measure the rotation of material frames $\{\mathbf{t}^{i},\mathbf{m}_{1}^{i},\mathbf{m}_{2}^{i}\}$ relative to the Bishop frame along $\mathbf{t}^{\textit{i}}$. $\mathbf{m}_{1}^{i},\mathbf{m}_{2}^{i}$ are defined as following
$\mathbf{m}_{1}^{i}=\cos\theta^{i}\cdot\textbf{u}^{i}+\sin\theta^{i}\cdot\textbf{v}^{i},\quad\mathbf{m}_{2}^{i}=-\sin\theta^{i}\cdot\textbf{u}^{i}+\cos\theta^{i}\cdot\textbf{v}^{i}.$ (13)
The material frames are used in (1) to obtain potential energy. An illustration of Bishop frame and material can be found in Figure 6. More relevant details can be found in [13, Section 4].
Appendix B Algorithmic Details
B.1 DNN Structure
@@
B.2 Verification of Improved Inextensiblity Enforcement with PBD
First, note that (9) applies only to successive vertices. This simplifies the conservation of momentum equation(12) so that it only involves successive vertices, as shown below:
$\textit{{M}}^{i}\cdot\Delta\hat{\textbf{x}}^{i}_{t+1}+\textit{{M}}^{i+1}\cdot\Delta\hat{\textbf{x}}^{i+1}_{t+1}=\textbf{0}$ (14)
$\textbf{r}^{i}\times(\textit{{M}}^{i}\cdot\Delta\hat{\textbf{x}}^{i}_{t+1})+\textbf{r}^{i+1}\times(\textit{{M}}^{i+1}\cdot\Delta\hat{\textbf{x}}^{i+1}_{t+1})=\textbf{0}$ (15)
The complete proposed solution, as detailed in Section 4.3, can be expressed as:
$\Delta\hat{\textbf{x}}^{i}_{t+1}=\frac{\textit{{M}}^{i+1}}{(||\textit{{M}}^{i}||_{2}+||\textit{{M}}^{i+1}||_{2})}\cdot C(\hat{\textbf{x}}^{i}_{t+1},\hat{\textbf{x}}^{i+1}_{t+1})\frac{\hat{\textbf{x}}^{i+1}_{t+1}-\hat{\textbf{x}}^{i}_{t+1}}{||\hat{\textbf{x}}^{i+1}_{t+1}-\hat{\textbf{x}}^{i}_{t+1}||_{2}}$ (16)
$\Delta\hat{\textbf{x}}^{i}_{t+1}=\frac{\textit{{M}}^{i}}{(||\textit{{M}}^{i}||_{2}+||\textit{{M}}^{i+1}||_{2})}\cdot C(\hat{\textbf{x}}^{i}_{t+1},\hat{\textbf{x}}^{i+1}_{t+1})\frac{-(\hat{\textbf{x}}^{i+1}_{t+1}-\hat{\textbf{x}}^{i}_{t+1})}{||\hat{\textbf{x}}^{i+1}_{t+1}-\hat{\textbf{x}}^{i}_{t+1}||_{2}}$ (17)
Expanding (14) with (16) and (17):
$\begin{split}&\textit{{M}}^{i}\cdot\Delta\hat{\textbf{x}}^{i}_{t+1}+\textit{{M}}^{i+1}\cdot\Delta\hat{\textbf{x}}^{i+1}_{t+1}=\frac{\textit{{M}}^{i}\cdot\textit{{M}}^{i+1}}{(||\textit{{M}}^{i}||_{2}+||\textit{{M}}^{i+1}||_{2})}\cdot C(\hat{\textbf{x}}^{i}_{t+1},\hat{\textbf{x}}^{i+1}_{t+1})\frac{\hat{\textbf{x}}^{i+1}_{t+1}-\hat{\textbf{x}}^{i}_{t+1}}{||\hat{\textbf{x}}^{i+1}_{t+1}-\hat{\textbf{x}}^{i}_{t+1}||_{2}}\\ &-\frac{\textit{{M}}^{i+1}\cdot\textit{{M}}^{i}}{(||\textit{{M}}^{i}||_{2}+||\textit{{M}}^{i+1}||_{2})}\cdot C(\hat{\textbf{x}}^{i}_{t+1},\hat{\textbf{x}}^{i+1}_{t+1})\frac{\hat{\textbf{x}}^{i+1}_{t+1}-\hat{\textbf{x}}^{i}_{t+1}}{||\hat{\textbf{x}}^{i+1}_{t+1}-\hat{\textbf{x}}^{i}_{t+1}||_{2}}\end{split}$ (18)
Since DLOs are linear and symmetric about their center axis, their mass matrices are also symmetric [43]. Consequently, $\textit{{M}}^{i+1}\cdot\textit{{M}}^{i}=\textit{{M}}^{i}\cdot\textit{{M}}^{i+1}$, allowing us to cancel both terms in (18) , resulting in zero linear momentum change. A key observation here is that (16) and (17) are collinear along the edge $i$. Therefore, $\Delta\hat{\textbf{x}}^{i}_{t+1}$ and $\Delta\hat{\textbf{x}}^{i+1}_{t+1}$ will not change orientation of edge $i$, naturally satisfies (15).
B.3 Summary and Discussion of Improved Inextensiblity Enforcement with PBD